[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Thu Aug 24, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Thu Aug 24 20:58:44 UTC 2017


HSU Shift Log for Thu Aug 24 2017 16:58:36 GMT-0400 (EDT)

Shift: H

Forecaster: R Berg

Technical Issues Summary:

ATCF issues:

N-AWIPS issues: Ridges on the Atlantic desk was freezing on me at one point  
early afternoon when I went to load new data. It eventually unfroze by  
itself.

AWIPS-2/NCP issues:

Other Techincal issues:

Miscellaneous issues:
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170824/6a23fff5/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list